OpenStudy (anonymous):

Find two non-negative numbers whose sum is 53 and such that the product of the two numbers is as large as possible. What is the global maximum?

zepdrix (zepdrix):

So we have 2 unknown numbers x and y which add up to 53, \[\Large x+y=53\] And they also have some product,\[\Large P=x\cdot y\]

zepdrix (zepdrix):

For this problem, we want to `maximize` the product. So we'll need to find `critical points` of the product function P. So we need \(\Large P'\) But before we go looking for the derivative, we want P in terms of one variable.

zepdrix (zepdrix):

We can use the x+y=53 to replace one of the variables in our P function.

zepdrix (zepdrix):

Understand what I mean by that? :o

OpenStudy (anonymous):

okaaay so far i think i understand. so if we choose to find x first then x= 53 - y

zepdrix (zepdrix):

good good. So what does our product function look like now?

OpenStudy (anonymous):

P = (53-y)(y)

zepdrix (zepdrix):

good good good.

OpenStudy (anonymous):

so 53y - y^2

zepdrix (zepdrix):

Ok cool. Now we want to take the derivative of our Product function with respect to y.

OpenStudy (anonymous):

i understand now! hehe it was a lot easier than i thought. thank you!

zepdrix (zepdrix):

yay \c:/